Show that
i. σ’x2 = σ’y2 = σ’z(2) = 1
ii. [σ’x , αx ] = 0 ,
[σ’x , αy ] = 2i αz &
[σ’x , αz] = -2i αy
1
Expert's answer
2020-07-06T17:01:40-0400

In order to prove this statements, we will use such notations:

"\sigma_x = \begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix},
\sigma_y = \begin{pmatrix}
 0 & -i \\
 i & 0
\end{pmatrix},
\sigma_z = \begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix}"

As we can see the first statement is right:

"\begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix}\begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix} = \begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix} = I"

"\begin{pmatrix}
 0 & -i \\
 i & 0
\end{pmatrix}\begin{pmatrix}
 0 & -i \\
 i & 0
\end{pmatrix} = \begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix} = I"

"\begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix}\begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix} = \begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix} = I"

The commutator of the two matrices is:

"\left[\sigma_x,\sigma_y\right] = \sigma_x\sigma_y - \sigma_y\sigma_x = \begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix}\begin{pmatrix}
 0 & -i \\
 i & 0
\end{pmatrix} - \begin{pmatrix}
 0 & -i \\
 i & 0
\end{pmatrix}\begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix} = 2i\begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix}"

So we have proved the second statement

Let's finish our work and prove the last statement:

"\left[\sigma_x,\sigma_z\right] = \sigma_x\sigma_z - \sigma_z\sigma_x = \begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix}\begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix} - \begin{pmatrix}
 1 & 0 \\
 0 & -1
\end{pmatrix}\begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix} = -2i\begin{pmatrix}
 0 & -i \\
 i & 0
\end{pmatrix}"